Reports of 21762
DGS (Test bench)
shoichi.oshino - 18:12 Friday 03 February 2023 (23828) Print this report
Preparation of Matlab on the workstation of test bench
To modify RT models on test bench system, I prepared Matlab for the workstation of test bench.
First I copied Matlab 2016b from k1nfs0.
After activating the license, I found that Matlab takes a long time to start.
I could not solve this problem, so I decided to start Matlab by specifying the license file directly.
I also copied k1export.m and startup.m files from the current k1boot.
When I opened the RT model file, I noticed that the link to the CDS parts was broken.
Previously CDS parts are installed in /opt/rtcds/rtscore, but in the new k1boot it is installed in /usr/share/advligorts.
So I cannot use the files via NFS.
Since advligorts can now be installed with apt, I will install this library directly in WS.
DGS (General)
satoru.ikeda - 18:07 Friday 03 February 2023 (23831) Print this report
Picomotor and Steppermotor Scripts and MEDM Changes

Overview
After starting the Picomotor and Steppermotor scripts, the EPICS channel is displayed very slowly, so the EPICS channel is always displayed and the communication part with the motor driver is changed to an on/off mechanism.

Operation

Please reopen sitemap first.
Otherwise, the old MEDM will be displayed.

In case of Picomotor

Press the blue OFF/ON button to start communication.
If it turns ON, the motor can be operated as before.

If communication fails, it turns OFF.
In this case, reboot with the rebooter and turn it ON in the same way.
Please turn it OFF after use.

In case of Steppermotor

Turn on the BIO and press the blue OFF/ON button to start communication.
If it turns ON, the motor can be operated as before.

If communication fails, it will turn OFF. In this case, turn BIO back on.
Please turn OFF after use.

Images attached to this report
Non-image files attached to this report
DGS (General)
satoru.ikeda - 17:58 Friday 03 February 2023 (23830) Print this report
Comment to Boot up k1script0 (23727)

Maintenance of k1script*

Changed IP address of k1script1's Pico network and connected k1script0 to Pico network.
DGS side is not changed.

k1script0: .250
k1script1: .250->.251

IPAddress

The address of the k1script in the authorized_keys file on the k1hwp* side of the HWP was also changed to reflect this change.

Non-image files attached to this comment
VIS (General)
satoru.ikeda - 17:49 Friday 03 February 2023 (23829) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

DGS Regular maintenance day(2/3)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO
  Check if the script can be started
  Checked that the limit switches installed are green.
 [Check result]
  PR0(PRM_GAS,PR3_GAS):NG
   => Network SW was abnormal. Recovered.

Release Sensor
 [Checked the Release Sensor.]
 Checked that Release Sensor is 0, 1: Seated.
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Confirmation]
  PingTest is performed.
 [Result]
  Only known

 NG:
  MCI,STM1,
  POS_POM,POM1
   => Confirmation only, no recovery operation was performed.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(IM,BF,STM)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,IMMT2,OMMT1,OMMT2,OSTM
  REFL_WFS,POP,POP2,POS,POS2,POP_WFS,AS_WFS
  POP_POM, MCF(MCO,MCI)
  PCAL(EX1,EX2,EY1,EY2):Not checked because it is ON only when used

Temperature controller
 [Checked]
 Is the temperature acquisition working properly?

 [Check result] 
  Is the temperature acquisition working properly?
  => No problem

Precision air Processer
 [Check Contents]
 Is the temperature acquisition working properly?

 [Result]
  No problem

FunctionGenerator
 (Confirmation)
 Is there a ping response?
 Are all EPICS channels alive?
  [sitemap]-[Commissioning]-[ALS FG]-[Xarm],[Yarm].
 [Check result]
 X and Y PLLs are working properly.

HWP control PC
 (Confirmation)
 Check if it is possible to login to the control PC.
 (Result) 
 Both k1hwp0, 1 and 2 are OK.
 

Images attached to this comment
Non-image files attached to this comment
VAC (Valves & Pumps)
nobuhiro.kimura - 15:25 Friday 03 February 2023 (23826) Print this report
Comment to Results of leak test for OMMT-OMC on 2nd/Feb./2023 (23802)

When:

3rd/Feb./2023

Workers:
N. Kimura
H. Yasui
 

Task: 
We performed a vacuum leak test for OMMT and OMC fter leak repair (https://klog.icrr.u-tokyo.ac.jp/osl/?r=23775) using a helium leak detector and helium gas. 

The tests were performed on the area on the AS area side (+X side of OMC) of the OMC and others.
The test was conducted with a test sensitivity of ~1 x 10^-12 Pam^3/s, confirming that there was no leakage > ~1 x 10^-11 Pam^3/s  .
Helium background during the test ranged aroud  ~1.55 x 10^-10 Pam^3/s. 
We conclud there are no leakeage on OMMT and OMC lager than ~1 x 10^-11 Pam^3/s, and can open the gate valve between SRM and OMMT.

VAC (Valves & Pumps)
nobuhiro.kimura - 15:16 Friday 03 February 2023 (23827) Print this report
Start-up of air compressor for gate valve drive

When:
3rd/Feb./2023
Workers:
N. Kimura
H. Yasui
 
Task: 
We started up the air compressor installed in the passage next to the computer room used to open and close the gate valve installed between SRM and OMMT
After startup, we inspected the piping connections for leaks and found a leak in the threaded part of the fitting.
(See attached photo)
To repair the leak, we replaced the seal at the threaded part.
Currently, the air compressor is in operation and compressed air for driving the gate valve is supplied to the gate valve.

Images attached to this report
VAC (Valves & Pumps)
nobuhiro.kimura - 14:52 Friday 03 February 2023 (23825) Print this report
Comment to Summaries of Ion Pump Activation (23668)

Abstract:

We checked the operational status of the activated ion pumps of IXV and IYV.
See below for comfirmed status. The activated ion pumps of IXV and IYV are in normal operation.

Workers:
N. Kimura
H. Yasui

Center area:
IXV (Position: Center 2F):
Status: Activated, evacuated by Ion Pump
11:26, 3/Feb./2023
The pressure displayed on the pressure gage: 4.4 x 10^-6 Pa
The pressure displayed on the Ion pump power supply: 2.4 x 10^-6 Pa
The anode voltage and current displayed on the Ion pump power supply: 6.2 kV and 363  x 10^-6 A

IYV (Position: Center 2F):
Status: Activated, evacuated by Ion Pump
11:27, 3/Feb./2023
The pressure displayed on the pressure gage :not confirmed
The pressure displayed on the Ion pump power supply: 2.5 x 10^-6 Pa
The anode voltage and current displayed on the Ion pump power supply: 6.1 kV and 372  x 10^-6 A

MIF (General)
naoatsu.hirata - 13:52 Friday 03 February 2023 (23824) Print this report
Windshield installation,PR2-POP

[Ushiba-san, Yokozawa-san, Takase-san, Hirata]

We tried to set windshield duct between PR2 flange and POP table wall. Unfortunately, the lenght is still little bit long. It can be installed by disassembling windshield itself, but it's annoying. So we decided to modify length, and try installing again next week.

MIF (ITF Control)
takafumi.ushiba - 13:37 Friday 03 February 2023 (23822) Print this report
Comment to PRC, MICH, and DARM filter modification for handover to OMC (23816)

Detail:

Currently, error signals of PRCL, MICH, and DARM was fractuated at low frequency in the current PRFPMI_RF_LOCKED state, so we need to modified control loops to supress the residual motion of each DoFs.
Figure 1, 2 and 3 show the new OLTF of PRCL, MICH, and DARM control, respectively.

Figure 4, 5, and 6 show the error signals of PRCL, MICH, DARM, when changing filters from old one to new ones.
T cursor chows the timing when engaging filters.
Error signals of PRL, MICH, and DARM becomes smaller and the loop seems stable enough.
So, I will try to switch TRAS signals to OMC DC for the next step.

Images attached to this comment
MIF (ASC)
tomotada.akutsu - 12:43 Friday 03 February 2023 (23821) Print this report
Comment to ADS of OMMT2 and OSTM for fine alignment to OMC (23815)

By the way, is this transmission power of OMC consistent? Maybe yes as follows, but please someone check it precisely:

  • IMC output around 2.2 W
  • PRM reduces it 1/10, so it becomes 220 mW.
  • Before reaching OMC, the beam passes BS twice, so 1/2*1/2 = 1/4, meaning 55 mW.
MIF (ASC)
takafumi.ushiba - 12:31 Friday 03 February 2023 (23820) Print this report
Comment to ADS of OMMT2 and OSTM for fine alignment to OMC (23815)

Detail:

First, I implemented OMMT2T centering loop to OMMT1, which is reported in klog23728.
Then, We tested ADSs of OMMT2 and OSTM for fine alignment to OMC.
Since the OMC QPD loop is currently stable (it was not stable during O3 due to glitch problems), we feedback ADS signals to the setpoint of QPD loops.

Followings are the detail of setting of ADSs:
Dither frequency: 15 Hz, 16 Hz, 17 Hz, and 18 Hz for OMMT2P, OMMT2Y, OSTMY, and OSTMP, respectively.
Dither amplitude: 200 cnts for pitch and 600 cnts for yaw.
INF filter: FM1 (BandPass filter)
OUTF filter: FM1 (LP1), FM2 (comb15), FM3 (comb1), FM9 (integrator)

Figure 1 shows the time series of ADS feedback and OMC output.
Transmission power increases from 34 to 50, and ADSs seem working well.
I implemnted the ADSs in INITIAL_ALIGNMENT guardian but maybe it is better to move it some other guardian like OMC_ASC.

Note:

During the work, OMC_ASC guardian moved OMMT1 and kicked the suspension because it was used for OMMT2T QPD centering.
Since OMMT1 doesn't use OMC alignment, I removed the request to OMMT1 from OMC_ASC guardian not to conflict the alignment control.

Also, we faced some bugs of OMC_ASC guardian.
Figure 2 shows the log around the bug.
It seems new thread is called withouut killing the existing thread, so it needs to be fixed.

Images attached to this comment
MIF (ASC)
tomotada.akutsu - 11:21 Friday 03 February 2023 (23818) Print this report
Comment to ADS of OMMT2 and OSTM for fine alignment to OMC (23815)

In the middle of the work, the OMC_ASC guardian threw an error on python-threading and die, and we spent a great deal of time working around it; not resolved yet but we just somehow "omit" it now. It appears to be due to a conflict among an already-existing thread and newly started one.

Although this happened today largely due to an unfortunate coincidence or other factors (like some unexpected lines in the OMC_LSC guardian, so we modified them then), nevertheless, it means that there is an error mixed in with the paths that could be taken, so one should do modify it.

VIS (IMM)
takaaki.yokozawa - 6:11 Friday 03 February 2023 (23819) Print this report
IMMT2 guardian stopped at aligning state
After finishing the X arm alignment with IR, I noticed that IMMT2 guardian stopped in aligning state.
The message is
Now waiting RMS < 10: TM_OLDAMP_P TM_OLDAMP_Y
So, temporary I added the number to opticalign
K1:VIS-IMMT2_TM_OPTICALIGN_P_OFFSET = -1300
K1:VIS-IMMT2_TM_OPTICALIGN_Y_OFFSET = -3600
Do we need the offload by pico motor?
MIF (General)
tomotada.akutsu - 0:19 Friday 03 February 2023 (23817) Print this report
Noise coherence check

Following 23811 and 23790, I checked coherence among DARM and NAB PDs in the wider frequency band when PRFPMI locked today (23816).

  • EXA NAB PD seems not working.
  • IYA NAB PD seems somehow unstable in high frequency. Time series signal might show glitchy feature, but I have no idea now.
  • Within this frequency range, the highest frequency having coherence seems no more than 10 Hz.

So far 2.8Hz? feature would have high coherence. This might be due to alignment flcutuation of certain mirrors might be converted to intra-cavity power or so. Further check with transmission power or TRX/Y or Oplevs might be helpful. Anyway it is out of the observation band.

Images attached to this report
MIF (ASC)
takafumi.ushiba - 23:14 Thursday 02 February 2023 (23815) Print this report
ADS of OMMT2 and OSTM for fine alignment to OMC

[Akutsu (remote), Ushiba]

Abstract:
We engaged ADSs to OMMT2 and OSTM for fine alignment to OMC.
It seems working well and transmitted power is increased by a factor of 1.3.

Detail will be posted tomorrow.

Note:
During this work, we reduced laser power as already repoted in klog23814.
During the work, we modified OMC_ASC and OMC_LSC guardian slightly.
It doesn't affect the OMC lock but if you find any porblems, please let us know.

Comments to this report:
tomotada.akutsu - 11:21 Friday 03 February 2023 (23818) Print this report

In the middle of the work, the OMC_ASC guardian threw an error on python-threading and die, and we spent a great deal of time working around it; not resolved yet but we just somehow "omit" it now. It appears to be due to a conflict among an already-existing thread and newly started one.

Although this happened today largely due to an unfortunate coincidence or other factors (like some unexpected lines in the OMC_LSC guardian, so we modified them then), nevertheless, it means that there is an error mixed in with the paths that could be taken, so one should do modify it.

takafumi.ushiba - 12:31 Friday 03 February 2023 (23820) Print this report

Detail:

First, I implemented OMMT2T centering loop to OMMT1, which is reported in klog23728.
Then, We tested ADSs of OMMT2 and OSTM for fine alignment to OMC.
Since the OMC QPD loop is currently stable (it was not stable during O3 due to glitch problems), we feedback ADS signals to the setpoint of QPD loops.

Followings are the detail of setting of ADSs:
Dither frequency: 15 Hz, 16 Hz, 17 Hz, and 18 Hz for OMMT2P, OMMT2Y, OSTMY, and OSTMP, respectively.
Dither amplitude: 200 cnts for pitch and 600 cnts for yaw.
INF filter: FM1 (BandPass filter)
OUTF filter: FM1 (LP1), FM2 (comb15), FM3 (comb1), FM9 (integrator)

Figure 1 shows the time series of ADS feedback and OMC output.
Transmission power increases from 34 to 50, and ADSs seem working well.
I implemnted the ADSs in INITIAL_ALIGNMENT guardian but maybe it is better to move it some other guardian like OMC_ASC.

Note:

During the work, OMC_ASC guardian moved OMMT1 and kicked the suspension because it was used for OMMT2T QPD centering.
Since OMMT1 doesn't use OMC alignment, I removed the request to OMMT1 from OMC_ASC guardian not to conflict the alignment control.

Also, we faced some bugs of OMC_ASC guardian.
Figure 2 shows the log around the bug.
It seems new thread is called withouut killing the existing thread, so it needs to be fixed.

Images attached to this comment
tomotada.akutsu - 12:43 Friday 03 February 2023 (23821) Print this report

By the way, is this transmission power of OMC consistent? Maybe yes as follows, but please someone check it precisely:

  • IMC output around 2.2 W
  • PRM reduces it 1/10, so it becomes 220 mW.
  • Before reaching OMC, the beam passes BS twice, so 1/2*1/2 = 1/4, meaning 55 mW.
MIF (ITF Control)
takafumi.ushiba - 23:13 Thursday 02 February 2023 (23816) Print this report
PRC, MICH, and DARM filter modification for handover to OMC

[Akutsu (remote), Ushiba]

Abstract:
We modified filter shape of PRCL, MICH, and DARM for handover TRAS signals to OMC.
The new filters seem working well, so we will try DC readout soon.

Detail will be posted tomorrow.

Comments to this report:
takafumi.ushiba - 13:37 Friday 03 February 2023 (23822) Print this report

Detail:

Currently, error signals of PRCL, MICH, and DARM was fractuated at low frequency in the current PRFPMI_RF_LOCKED state, so we need to modified control loops to supress the residual motion of each DoFs.
Figure 1, 2 and 3 show the new OLTF of PRCL, MICH, and DARM control, respectively.

Figure 4, 5, and 6 show the error signals of PRCL, MICH, DARM, when changing filters from old one to new ones.
T cursor chows the timing when engaging filters.
Error signals of PRL, MICH, and DARM becomes smaller and the loop seems stable enough.
So, I will try to switch TRAS signals to OMC DC for the next step.

Images attached to this comment
IOO (IMC)
takafumi.ushiba - 23:07 Thursday 02 February 2023 (23814) Print this report
Power reduction of PSL output

Since the OMC output power is close to 60 mW when OMC alignment is good, I reduced PSL output power by rotating PSL HWP.
Now, PSL output  power is about 2.2 W while it was 2.8 W before rotating HWP.
Since signals used for LSC except for CARM is normalized by IMC trans power, it doesn't affect the lock of FPMI (though CARM gain was reduced slightly but robust enough).

I'm not sure single arm lock is stable now (I hope it is also fine though), so if single arm lock is unstable please increase the laser power or adjust the gain of CARM CMS.

VIS (EX)
Kohei Mitsuhashi - 19:08 Thursday 02 February 2023 (23812) Print this report
tried to match the value of the power spectra
This log report covers two aspects:
1) I tried to match the values of power spectra measured by the accelerometers and the geophones (nemed ETMX_IP_ACCINF and ETMX_IP_FLDACC).
2) I tried to match the values of power spectra measured by the accelerometers and the geophones (nemed ETMX_IP_ACCBLEND_ACCL, T, Y and ETMX_IP_FLDACC_FLDACCL, T, Y)


I tried to match the values of power spectra measured by the accelerometer and the geophones (nemed ETMX_IP_ACCINF and ETMX_IP_FLDACC).
To do this, I changed the servo filter named ETMX IP FLDACCINF FILTER.
The picture of results was attached.

I tried to match the values of power spectra measured by the accelerometer and the geophones (nemed ETMX_IP_ACCBLEND_ACCL, T, Y and ETMX_IP_FLDACC_FLDACCL, T, Y).
To do this, I changed the matrix named ETMX IP FLDACCALIGN MATRIX.
The picture of results was attached.
Images attached to this report
MIF (General)
tomotada.akutsu - 18:06 Thursday 02 February 2023 (23811) Print this report
NAB PD at EXA not turned on?

Following 23790, I checked some NAB PDs and found that one at EXA seemed strange. Are they turned on???

Images attached to this report
DetChar (General)
shoichi.oshino - 17:08 Thursday 02 February 2023 (23810) Print this report
Add high pass filter to the strain channel on SummaryPage
Yamamoto-san mentioned that 10-70Hz of DARM sensitivity may be polluted by numerical error.
This is because the ASD of the strain channel differs greatly between low and high frequencies.
This difference is even greater at PRFPMI and needs to handle.

I compared the strain plotted in the following three settings.
1. FFT length 4s
2. FFT length 60s
3. FFT length 4s + high pass filter 10Hz

The data used are the following times:
LSC_LOCK guardian state : PRFPMI_RF_LOCKED
GPS time : 1359200000 - 1359200600

The attached files are the strain of each setting.
2 and 3 improve the strain by a few tens of Hz.
Therefore, I choose setting 3 for SummaryPage.
Images attached to this report
MIF (ITF Control)
takafumi.ushiba - 16:03 Thursday 02 February 2023 (23809) Print this report
Performance test of PNCs

Abstract:
I measured spectra of ALS signals when FPMI was locked and compared the spectra with and without PNCs.
large peaks at low frequency in ALS_DARM signals are completely disappeared while ALS_CARM still has a peak at 0.46 Hz.
Since the peak in ALS_CARM and PR2 TML OpLev has a large coherence, it comes from PR2 motion.

Detail:
Figure 1 shows the spectra of ALS signals when FPMI was locked.
Reference are the signals without PNC and current values are the signals with PNC.
Large peaks at 0.38 Hz and 0.43 Hz are completely disappeared from ALS signals by engaging PNC.

However, ALS_CARM signals still have a peak at 0.46 Hz while ALS_DARM signals don't have.
The fact that ALS_DARM doesn't have a peak implies that the source of this peak is coming from IR path.
So, I checked the coherence of ALS_CARM signals and PR2 TML OpLev (right bottom graph in the fig1).
There is a large coherence around 0.46 Hz, so we ned to damp 0.46 Hz longitudinal mode of PR2 to improve further.

Images attached to this report
DGS (Test bench)
shoichi.oshino - 15:49 Thursday 02 February 2023 (23808) Print this report
RT model build
I tried to build RT model on k1boot of the test bench.
First I created /opt/rtcds/userapps/release directory to put model files.
Also created /opt/rtcds/userapps/release/etc/userapps-user-env.sh file to set the path of RT models to the RCG_LIB_PATH parameter.
I found that the directory structure was different from the current KAGRA DGS at this time.
This is because of my misunderstanding of facter and hiera so I modified "site" parameter to KAMIOKA.
The build command now works, but an error has occurred so I will fix this.
VIS (OMM)
ryutaro.takahashi - 15:24 Thursday 02 February 2023 (23807) Print this report
Comment to Health check of OMMT1/2/OSTM suspensions (22937)

This is a report of the health check in the OMMT1/2/OSTM suspensions after the pump-down.
The measured transfer functions look healthy. There are some deviations from the optimal position of OSEM.

OMMT1 H3: horizontal
OMMT1 H4: axial, horizontal
OMMT2 H1: axial, horizontal
OMMT2 H3: axial
OSTM H1: axial
OSTM H2: axial, horizontal

Images attached to this comment
VAC (Valves & Pumps)
nobuhiro.kimura - 15:18 Thursday 02 February 2023 (23802) Print this report
Results of leak test for OMMT-OMC on 2nd/Feb./2023

When:

2nd/Feb./2023

Workers:
N. Kimura
H. Yasui
M. Takahashi 
 

Task: 
We performed a vacuum leak test for OMMT and OMC fter leak repair (https://klog.icrr.u-tokyo.ac.jp/osl/?r=23775) using a helium leak detector and helium gas. 

The tests were performed on all flange connections except the top flange of the OMMT and OMC and the area on the ASC side of the OMC.
The test was conducted with a test sensitivity of ~1 x 10^-12 Pam^3/s, confirming that there was no leakage > ~1 x 10^-11 Pam^3/s  .
Helium background during the test ranged from 8.5x 10^-10 Pam^3/s to 1 x 10^-10 Pam^3/s. 
Leakage testing of the flanges in the area on the ASC side is scheduled for 3rd/Feb..

Images attached to this report
Comments to this report:
nobuhiro.kimura - 15:25 Friday 03 February 2023 (23826) Print this report

When:

3rd/Feb./2023

Workers:
N. Kimura
H. Yasui
 

Task: 
We performed a vacuum leak test for OMMT and OMC fter leak repair (https://klog.icrr.u-tokyo.ac.jp/osl/?r=23775) using a helium leak detector and helium gas. 

The tests were performed on the area on the AS area side (+X side of OMC) of the OMC and others.
The test was conducted with a test sensitivity of ~1 x 10^-12 Pam^3/s, confirming that there was no leakage > ~1 x 10^-11 Pam^3/s  .
Helium background during the test ranged aroud  ~1.55 x 10^-10 Pam^3/s. 
We conclud there are no leakeage on OMMT and OMC lager than ~1 x 10^-11 Pam^3/s, and can open the gate valve between SRM and OMMT.

VAC (Valves & Pumps)
nobuhiro.kimura - 15:14 Thursday 02 February 2023 (23806) Print this report
Comment to Results of leak test for X-15 vacuum pump unit (23804)

When:
1/Feb./2023
From 9:50 to 9:53
 
Workers:
N. Kimura
H. Yasui
 
Task:
We connected activated Ion pump of X-15  to the X-arm duct. 
Status of the Ion pump and gate valves were as follows;
9:50 
Ion pump:
The pressure displayed on the pressure gage: 7.1 x 10^-7 Pa.
Pressure displayed on the Ion pump power supply: 5.4 x 10^-7 Pa
The anode voltage and current displayed on the Ion pump power supply : 5.9 kV and 78 x 10^-6 A
GV between X-arm duct and the pump unit: CLOSE
GV between Ion pump and TMP: OPEN
 
9:52  
GV between Ion pump and TMP: CLOSE
The pressure displayed on the pressure gage: 7.8 x 10^-7 Pa.
9:53  
GV between X-arm duct and the pump unit: OPEN
The pressure displayed on the pressure gage: 1.3 x 10^-6 Pa.
TMP turned off

Search Help
×

Warning

×